# Holographic drag force with translational symmetry breaking

### v1

@@ 1 Introduction
The holographic drag force formalism is explained in [3, 4, 5] as the momentum rate flowing down to the probe string is interpreted as the drag force exerted. In the gravitational side, the presence of probe branes in the AdS bulk breaks conformal symmetry and sets the energy scales so leads to corrections in $AdS_{5}$. Drag force was studied using Gubser’s proposal in the context of non-conformal models in [6, 7, 8, 9, 10].
@@
In this study, we investigate calculation of the drag force in systems where translational symmetry is not preserved. We employ holographic models that address this symmetry breaking. Although these models maintain a symmetric background metric, the symmetry is interrupted by introducing either a mass for the graviton or other fields. For example, Vegh’s model introduces mass terms for gravitons that interfere with the bulk diffeomorphisms, which correspond to translational symmetry in holography [20, 16]. Similarly, Andrade and Withers’ approach uses massless scalar fields that exhibit a linear relationship with a spatial coordinate in the field theory [21]. A model with a massless two-form field that aligns linearly with a spatial coordinate in a conformal field theory (CFT) was introduced in [22]. A comprehensive framework for translational symmetry breaking has been examined in [23], utilizing holography. This framework considers models of holographic massive gravity and manipulates translational symmetry, either explicitly or spontaneously, through a variable parameter. The dynamics of the drag force in a context of massive gravity have been previously explored in [24].
@@ 2 Translational symmetry breaking model
where $i=1,2,3$ are orthogonal spatial boundary coordinates, $a$ labels the $d-1$ spatial $x^{a}$ directions, $I$ is an internal index that labels the $d-1$ scalar fields and $\alpha_{I\,a}$ are real arbitrary constants. The ansatz (2) leads to the solution,
@@
As our current work aims to work on an asymptomatic $AdS_{5}$, let’s assume $d=4$ in the model (1) which is written as,
@@
respectively. Plugging (11) into (10) provides,
@@
and imposing the null energy condition on the Ricci tensor of the metric function in (2) gives,
@@
Therefore, in order to perform the calculations in the following sections, conditions (14) and (15) must be met.
@@ 3 Drag force in a TSB system
where $\mathcal{O}(t)$ are all terms that vanish at late time, from this point forward we ignore them. From (9) and (3) we find the Lagrangy as,
@@
and from (19) the energy-momentum current is written as,
@@
where denotes the derivation with respect to $r$ in this case. From (20) one can obtain,
@@
The right side of (21) needs to be real in order to prevent an imaginary string ansatz. Thus there should be a common root between the nominator and denominator. To apply this condition first we find the root of nominator as,
@@
where $r_{c}$ is the root of (22). Following this, we arrive at,
@@
As previously stated, the denominator of (21) must be satisfied by the above $r_{c}$ (3) as,
@@
where $r_{c}$ is written as (3) and $\sqrt{f(r_{c})}$ is found from (27). Note that the drag force’s negative sign originates from the force vector’s direction, so in order to study $F$ we ignore the --, also take the $12\pi\alpha^{\prime}=$ 1. We continue by examining the drag force’s behavior in plots. Keep in mind that we must verify constraints (14) and (15) before selecting different parameter values. Since (15) is obviously satisfied for all values, we only concentrate on (14).
@@
Figure 1: Drag force versus speed $v$ for a) different values of $\alpha$ and $\mu=1$, b) different values of $\mu$ and $\alpha=1$.
@@
Figure 2: Drag force a) versus chemical potential $\mu$, b) versus $\alpha$ for different values of $v$.
@@ 4 Drag force at zero temperature in a TSB system
Examine constraint (14) at precisely $T=0$, we may therefore discuss $F_{drag}$ in (29) by determining different values of $(\alpha,\mu)$ at some fixed $r_{0}$.
@@
Figure 3: Drag force at zero temperature versus $v$ for a) different values of $\mu$ , b) different values of $\alpha$.
The drag force on a semi-classical string object moving at zero temperature is depicted in Figure 3. Plot (a) indicates that increasing chemical potential $\mu$ at zero temperature increases drag force, whereas plot (b) demonstrates that increasing $\alpha$ has the opposite effect. We find the different effects of $\alpha$ in a zero temperature system and a thermal medium by comparing with figure 1.
Upon holding the system’s parameters constant and varying the temperature, it is observed that the heavy particle is subjected to an increased drag force when the temperature is zero. In field theory, the observation that entropy density remains non-zero even at zero temperature indicates the existence of states to which a point particle can dissipate energy and momentum. This suggests that despite the absence of thermal energy, there are other mechanisms or degrees of freedom within the system that allow for the transfer of energy and momentum from the particle.
@@ 5 Diffusion Coefficient
Therefore from (30) the diffusion coefficient is written as,
@@
Like we did for the drag force in the previous section, we examine how coefficient $D$ behaves in relation to various variables in the plots that follow.
@@ 6 Fluctuations on the string in translational breaking background
Heavy quarks out of equilibrium exhibit Brownian-like motion under the influence of a stochastic force, denoted as $\zeta(t)$. These dynamics produce observables that are essential to comprehending the behavior of the quark gluon plasma. One can find a thorough synopsis of the physical principles in [26]. In [27, 28, 29] an extensive research has been done on Langevin coefficients within the gauge/gravity duality. Also a universal methodology applicable to a broad spectrum of gravity theories, utilizing the membrane paradigm, was introduced in [30].
In this section, we study the relativistic Langevin coefficients in the translational symmetry background (2) using holography. Our main goal is to study these coefficients. It was argued in [31] that there is a universal inequality between transverse and longitudinal Langevin coefficients, and it only violates in an anisotropic background. We investigate its violation in the background (2). Here, we analyze the stochastic forces exerted on a quark that is in motion at a uniform speed. These forces arise due to the existence of a horizon on the string world-sheet, $r_{c}$. The Browniaan’s motion of the quark can be modeled by the generalized Langevin equations. These equations incorporate Langevin coefficients that are proportional to the temperature of the string world-sheet, reflecting the thermal context of the motion [31]. One finds the ratio of longitudinal to transverse Langevin coefficients as,

### v2

@@ 1 Introduction
The holographic drag force formalism is explained in [3, 4, 5] as the momentum rate flowing down to the probe string is interpreted as the drag force exerted. In the gravitational side, the presence of probe branes in the AdS bulk breaks conformal symmetry and sets the energy scales so leads to corrections in $AdS_{5}$. Drag force was studied using Gubser’s proposal in the context of non-conformal models in [6, 7].
@@
In this study, we investigate calculation of the drag force in systems where translational symmetry is not preserved. We employ holographic models that address this symmetry breaking. Although these models maintain a symmetric background metric, the symmetry is interrupted by introducing either a mass for the graviton or other fields. For example, Vegh’s model introduces mass terms for gravitons that interfere with the bulk diffeomorphisms, which correspond to translational symmetry in holography [17, 13]. Similarly, Andrade and Withers’ approach uses massless scalar fields that exhibit a linear relationship with a spatial coordinate in the field theory [18]. A model with a massless two-form field that aligns linearly with a spatial coordinate in a conformal field theory (CFT) was introduced in [19]. A comprehensive framework for translational symmetry breaking has been examined in [20], utilizing holography. This framework considers models of holographic massive gravity and manipulates translational symmetry, either explicitly or spontaneously, through a variable parameter.
@@ 2 Translational symmetry breaking model
where $a$ labels the $d-1$ spatial $x^{a}$ directions, $I$ is an internal index that labels the $d-1$ scalar fields and $\alpha_{I\,a}$ are real arbitrary constants. The ansatz Eq.(2) leads to the solution,
@@
As our current work aims to work on an asymptomatic $AdS_{5}$, let’s assume $d=4$ in the model Eq.(1) which is written as,
@@
respectively. As demonstrated by Eq.(12), since the temperature varies with $\alpha$ and $\mu$, it is not fixed, (therefore, when examining the drag force in the next section, we will make the quantities dimensionless).
Plugging Eq.(11) into Eq.(10) provides,
@@
and imposing the null energy condition on the Ricci tensor of the metric function in Eq.(2) gives,
@@
Therefore, in order to perform the calculations in the following sections, conditions Eq.(14) and Eq.(15) must be met.
@@ 3 Drag force in a TSB system
When a quark moves through a hot medium it quenches strongly. Drag force is related to the energy loss of heavy quark, based on the interaction between the quark and the medium. From string theory point of view the energy loss of the quark is understood as the momentum flow along a moving classical string into the horizon. In this conjecture the moving quark is mapped onto a probe string in the AdS space. So instead of studying the motion of quark in a strongly coupled system one can simply consider the classical dynamics of a string on the gravitational side. Since $\mathcal{N}=4$ SYM theory can capture most of the dynamics of QGP after a critical temperature, AdS/QCD approximate the phenomenology of QCD by replacing it with $\mathcal{N}=4$ SYM theory. According to the holographic drag force formalism momentum rate flowing down to the probe string is interpreted as the drag force and the quark is prescribed to move on the boundary of $AdS_{5}$. The bottom-up approach starts with a five-dimensional effective field theory somehow motivated by string theory and tries to fit it to QCD as much as possible (as examples see [21, 22, 23]), also in [24] transport properties of Quark-Gluon Plasma with a machine-learning assisted holographic approach is studied. In the linear-axion model, momentum relaxation is achieved by introducing a linear axion field, leading to a drag force that is proportional to the velocity of the probe particle [25]. This results in a linear drag force behavior. The model exhibits a transition between coherent and incoherent metal phases, which affects the drag force. Additionally, the longitudinal shear viscosity in this model can violate the Kovtun-Son-Starinets (KSS) bound, indicating strong anisotropic effects. Similarly, in the massive gravity model, the introduction of massive gravitons breaks the diffeomorphism invariance, resulting in momentum dissipation and a non-zero drag force even at zero temperature [26]. The drag force in massive gravity models is often linked to universal conductivity properties, with the presence of massive gravitons leading to a finite DC conductivity that affects the drag force behavior. Like the linear-axion model, the massive gravity model also shows phase transitions that impact the drag force, which can be studied through the behavior of the probe particle in the holographic setup. These features highlight some universal mechanisms, such as the role of momentum dissipation on drag force behavior.
@@
where $\mathcal{O}(t)$ are all terms that vanish at late time, from this point forward we ignore them. From Eq.(9) and Eq.(3) we find the Lagrangian as,
@@
and from Eq.(19) the energy-momentum current is written as,
@@
where denotes the derivation with respect to $r$ in this case. From Eq.(20) one can obtain,
@@
The right side of Eq.(21) needs to be real in order to prevent an imaginary string ansatz. Thus there should be a common root between the numerator and denominator. To apply this condition first we find the root of numerator as,
@@
where $r_{c}$ is the root of Eq.(22). We define the dimensionless variables $\tilde{\alpha}=\frac{\alpha}{2\pi T}$, $\tilde{\mu}=\frac{\mu}{2\pi T}$, $\tilde{r}=\frac{r}{2\pi T}$, $\tilde{r}_{0}=\frac{r_{0}}{2\pi T}$ in order to solve the above equation. The Eq.(22) is then expanded and written using the new dimensionless variables. We can eliminate the tilde sign in the following equation for simplicity’s sake.
@@
As previously stated, the denominator of Eq.(21) must be satisfied by the above $r_{c}$ Eq.(3) as,
@@
where $r_{c}$ is written as Eq.(3) and $\sqrt{f(r_{c})}$ is found from Eq.(27). Note that the drag force’s negative sign originates from the force vector’s direction, so in order to study $F$ we ignore the --, also take the $12\pi\alpha^{\prime}=$ 1. We continue by examining the drag force’s behavior in plots. Keep in mind that we must verify constraints Eq.(14) and Eq.(15) before selecting different parameter values. Since Eq.(15) is obviously satisfied for all values, we only concentrate on Eq.(14).
@@
Figure 1: Drag force versus speed $v$ for a) different values of $\alpha$ and $\mu=1$, b) different values of $\mu$ and $\alpha=1$. Plots are rescaled, but for the sake of simplicity, the quantities’ tilde sign is removed.
@@
Figure 2: Drag force a) versus chemical potential $\mu$, b) versus $\alpha$ for different values of $v$. Plots are rescaled, but for the sake of simplicity, the quantities’ tilde sign is removed.
@@ 4 Drag force at zero temperature in a TSB system
Examine constraint Eq.(14) at precisely $T=0$, we may therefore discuss $F_{drag}$ in Eq.(29) by determining different values of $(\alpha,\mu)$ at some fixed $r_{0}$.
@@
Figure 3: Drag force at zero temperature versus $v$ for a) different values of $\alpha$, b) different values of $\mu$.
The drag force on a semi-classical string object moving at zero temperature is depicted in Figure 3. Plot (a) demonstrates that increasing $\alpha$ at zero temperature decreases the drag force whereas plot (b) indicates that increasing chemical potential $\mu$ has the opposite effect. We find the different effects of $\alpha$ in a zero temperature system and a thermal medium by comparing with figure 1.
Upon holding the system’s parameters constant and varying the temperature, it is observed that the heavy particle is subjected to an increased drag force when the temperature is zero. In terms of field theory, we observe that the entropy density remains non-zero even at zero temperature. This implies the existence of states to which a point particle can lose energy and momentum. Essentially, the system retains a non-zero density of states, allowing for the dissipation of momentum and energy by the point particle, even in the absence of thermal excitation. From a physical perspective, this non-zero entropy density at zero temperature indicates that there is a sufficient density of states available for a string to radiate momentum. This means that, despite the temperature being zero, the string can still interact with these states, losing momentum in the process. The presence of these states is crucial as it ensures that the string has a medium to transfer its momentum to, thereby maintaining the dynamics of the system even in a zero-temperature environment.
Examining the behavior of the drag force versus temperature would be helpful, given the intriguing findings of studying drag force at finite temperature and zero temperature. Keep in mind that rescaling has no point on this computation. Thus, starting from (12) we write in this manner,
$T-\frac{1}{4\pi}\Big{(}4r_{0}-\frac{\alpha^{2}}{2\,r_{0}}-\frac{2\,\mu^{2}}{3r_{0}}\Big{)}=0,$ (30)
subsequently $r_{0}$ as a function of $T$ will be obtained by solving (30) for $r_{0}$. The drag force is expressed in terms of $\alpha$, $\mu$, and $T$ when we substitute $r_{0}$ in terms of T in (29).
Refer to captionRefer to $\end{array}$
Figure 4: Drag force versus T for arbitrary fixed values of $\alpha$ and $\mu$ and different values of $v$.
The drag force versus temperature for various values of $v$ is displayed in figure 4. It is demonstrated that the drag force has an almost monotonous behavior with respect to temperature when $\alpha$ and $\mu$ are fixed. On the other hand, when increasing velocity is approaching, this behavior may gradually change. When the velocity is high enough, as in the highest $v$ plot, the significant difference between the drag force at low temperature and high temperature is clearly expressed, even though the $F$ appears to be a constant function at the lowest $v$ plot.
(a)Refer to caption (b)Refer to caption
Figure 5: Drag force versus temperature for a) different values of $\alpha$ with arbitrary fixed values of $v$ and $\mu$ and b) different values of $\mu$ with arbitrary fixed values of $v$ and $\alpha$
Figure 5 plots also show F versus T in the following scenarios: a) $\mu$ and v are fixed but $\alpha$ is increasing; b) $\alpha$ and v are fixed but $\mu$ is increasing. Plot (a) demonstrates that the drag force develops a maximum value when $\alpha$ is increasing. This peak is apparent at low temperatures when $\alpha$ reaches a sufficient value, and at higher temperatures, the drag force plots with various $\alpha$ values are converging. Plot b) shows that, the drag force increases with increasing chemical potential as we knew from previous sections; however, there is no maximum $F$ when $\mu$ is increasing.
Backreactions typically become significant at extremely low temperatures, which suggests that the probe limit of the probe string approach may encounter issues at zero temperature. In such conditions, the interactions between the probe string and the background cannot be ignored, as they can lead to substantial deviations from the expected behavior. This implies that the simplifications assumed in the probe limit might break down, necessitating a more comprehensive treatment that accounts for the backreactions to accurately describe the system’s dynamics at zero temperature. Consequently, the reliability of the probe string approach in these extreme conditions is questionable, highlighting the need for alternative methods or adjustments to the existing framework to ensure accurate modeling.
@@ 5 Diffusion Coefficient
Therefore from Eq.(31) the diffusion coefficient is written as,
@@
Where $F$ is written as Eq.(29). Like we did for the drag force in the previous section, we examine how coefficient $D$ behaves in relation to various (dimensionless) variables in the plots that follow.
@@ 6 Fluctuations on the string in translational breaking background
Heavy quarks out of equilibrium exhibit Brownian-like motion under the influence of a stochastic force, denoted as $\xi(t)$. These dynamics produce observables that are essential to comprehending the behavior of the quark gluon plasma. One can find a thorough synopsis of the physical principles in [28]. In [29, 30, 31] an extensive research has been done on Langevin coefficients within the gauge/gravity duality. Also a universal methodology applicable to a broad spectrum of gravity theories, utilizing the membrane paradigm, was introduced in [32].
A quark moving with a constant velocity exhibits dynamics similar to Brownian motion. Its motion can be described using generalized Langevin equations, which incorporate the components of real-time correlation functions for the time-dependent $\xi(t)$. Assuming that, over long periods, the time-correlation functions are proportional to delta functions, the Langevin equations become local, and the diffusion coefficients remain constant. In [33], it was argued that a universal inequality exists between transverse $\kappa_{T}$ and longitudinal $\kappa_{L}$ Langevin coefficients in generic isotropic backgrounds. It was shown that this inequality is only violated in anisotropic backgrounds, leading to negative excess noise in the system. In this work, we investigate the violation of this inequality in the background described by (2). First we explain that the Browniaan’s motion of the quark can be modeled by the generalized Langevin equations. Two-point correlators for transverse and longitudinal to the direction of motion are given by the following equations
$\xi_{T}(t)\xi_{T}(t^{\prime})\right>=\kappa_{T}\delta(t-t^{\prime}),\,\,\,\,\,\left<\xi_{L}(t)\xi_{L}(t^{\prime})\right>=\kappa_{L}\delta(t-t^{\prime}),$ (34)
Where $\kappa_{L}$ and $\kappa_{R}$ are related to anti-symmetric retarded correlator $G_{R}$ as
$\kappa_{T}=-2T_{ws}\lim_{\omega\rightarrow 0}\frac{ImG_{R}^{T}(\omega)}{\omega},\,\,\,\,\,\,\kappa_{L}=-2T_{ws}\lim_{\omega\rightarrow 0}\frac{ImG_{R}^{L}(\omega)}{\omega}$ (35)
In these relations the world sheet temperature is denoted by $T_{ws}$.
Next, we study the relativistic Langevin coefficients in the translational symmetry background Eq.(2) using holography. Our main goal is to study these coefficients. Here, we analyze the stochastic forces exerted on a quark that is in motion at a uniform speed. These forces arise due to the existence of a horizon on the string world-sheet, $r_{c}$. These equations incorporate Langevin coefficients that are proportional to $T_{ws}$, reflecting the thermal context of the motion [33]. One finds the ratio of longitudinal to transverse Langevin coefficients as,